OpenStudy (anonymous):

Multiple choice Given f(x) = x^2 + 3 and g (x) = x+5/x Find (g o f) (4) 15 -15 14/19 24/19

OpenStudy (anonymous):

(g o f)(4) = g(f(4))

OpenStudy (anonymous):

First, find f(4). Then, substitute the value of f(4) into g(x)

OpenStudy (anonymous):

Im confused (gof) (-4) = g (f(-4

OpenStudy (anonymous):

would 15

OpenStudy (anonymous):

f(x) = x^2 + 3 g(x) = (x + 5) / x f(4) = 4^2 + 3 = 19 g(f(4)) = g(19) = (19 + 5) / 19 = 24 / 19